# Solving quantum-inspired dynamics on quantum and classical annealers

### v1

@@ Abstract
We propose a practical benchmarking suite inspired by physical dynamics to challenge both quantum and classical computers. Using a parallel in time encoding, we convert the real-time propagator of an $n$-qubit, possibly non-Hermitian, Hamiltonian into a quadratic-unconstrained binary optimisation (QUBO) problem. The resulting QUBO instances are executed on D-Wave quantum annealers as well as using two classical solvers, Simulated Annealing and VeloxQ, a state-of-the-art classical heuristic solver. This enables a direct comparison. To stress-test the workflow, we use eight representative models, divided into three groups: (i) single-qubit rotations, (ii) multi-qubit entangling gates (Bell, GHZ, cluster), and (iii) PT-symmetric, parity-conserving and other non-Hermitian generators. Across this diverse suite we track the success probability and time to solution, which are well established measures in the realm of heuristic combinatorial optimisation. Our results show that D-Wave Advantage2 consistently surpasses its predecessor, while VeloxQ presently retains the overall lead, reflecting the maturity of classical optimisers. We highlight the rapid progress of analog quantum optimisation, and suggest a clear trajectory toward quantum competitive dynamics simulation, by establishing the parallel in time QUBO framework as a versatile test-bed for tracking and evaluating that progress.
@@ II.1 Quantum Annealing
governed by QPU dependent anneal schedules $A(s)$ and $B(s)$ (cf. Fig. (LABEL:anneal_schedule)). If the evolution is sufficiently slow, according to the adiabatic theorem [27], the system remains in its instantaneous ground state, and the final state gives the solution to the original problem.
D-Wave’s superconducting quantum annealers are now applied across a spectrum of physics-driven tasks. Experiments on a 5000-qubit processor reproduced quantum critical dynamics in 3D spin glasses, extracting critical exponents beyond the reach of classical simulation [28]. In molecular science, an annealer-based eigensolver has recovered excited-state spectra for small molecules, showing that electronic-structure problems can be recast as Ising models amenable to quantum annealing [29]. Hybrid quantum-classical workflows have also tackled real-time traffic-flow optimisation, demonstrating measurable congestion reduction on realistic road networks [30, 31, 32, 33, 34]. In computer science, an adiabatic implementation of Simon’s algorithm has been recently proposed and solved on D-Wave quantum annealers [35]. Furthermore, the D-Wave machine has been used as a platform for thermodynamic experiments to validate quantum fluctuation theorems and simulate quantum thermal machines [36, 37, 38]. Progress notwithstanding, widespread deployment is limited by sparse qubit connectivity, which imposes costly minor-embeddings, and by analogue control errors and thermal noise that dilute prospective speed-ups [39, 40, 41, 42].
@@ II.2 VeloxQ
VeloxQ is a fast and efficient, physics-inspired solver based on the theory of dynamical systems, designed for solving QUBO problems, which are crucial for various real-world optimisation tasks [43]. In a recent work, it has been extensively benchmarked against many state-of-the-art QUBO solvers, D-Wave quantum annealers among them [16]. Unlike them (or other physics-inspired optimisation approaches such as gate-based quantum computers), VeloxQ does not rely on technological advancements in hardware to reach its full potential, as it leverages conventional computing technology (such as GPUs). While it is a classical solver, it emulates the probabilistic nature of quantum annealers by producing a low-energy spectrum of solutions. By default, the solver operates with a carefully selected set of parameters that include instance-dependent tuning. However, it also offers flexibility, allowing users to adjust parameters to control the trade-off between solution quality and runtime based on problem knowledge or prior runs. Crucially, it offers stable and repeatable results, which do not suffer from unstable hardware.
VeloxQ distinguishes itself from D-Wave quantum annealers in several significant ways. A key difference lies in their handling of problem topology: D-Wave’s quantum processing units (QPUs) have a fixed spin connectivity structure, and necessitate “minor embedding” for most real-world problems due to their specific topologies, which introduces substantial overhead. In contrast, VeloxQ natively handles all problem topologies without requiring such embedding. Furthermore, VeloxQ exhibits superior scalability, capable of tackling large-scale optimisation problems with up to $2\times 10^{8}$ sparsely connected variables, a feat currently intractable for D-Wave and its hybrid solutions, with estimates suggesting decades before quantum annealers could reach similar capabilities [16].
@@ II.4 Problem Instances
We examine entangled and non-entangled quantum system, as well as non-Hermitian Hamiltonians that model quantum dynamics such as $\mathcal{P}\mathcal{T}$-symmetric qubits. To perform an accurate performance comparison between classical and quantum solvers, we select dynamical systems that can be precisely modelled as QUBO even with limited precision. Some of these dynamical systems produce QUBO instances that are native to the Pegasus graph, others require an embedding.
@@
Besides Hermitian operators, which are usually used as observables in quantum theory, quasi-Hermitian Hamiltonians allow to model unitary evolution as well. We simulate the behavior of a specific type of quasi-Hermitian system, the so-called $\mathcal{P}\mathcal{T}$-symmetric Hamiltonians. $\mathcal{P}\mathcal{T}$-symmetric systems remain invariant under the combined parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$) operator. Both of these transformations are hermitian and independent of each other, with $\mathcal{P},\mathcal{T}]=0 [$45]. Notably, systems with unbroken $\mathcal{P}\mathcal{T}$ symmetry have a real eigenspectrum and exhibit quasi-Hermitian behavior. A system is unbroken if the eigenstates of $H$ correspond to the eigenstates of the $\mathcal{P}\mathcal{T}$ operator. We evaluate the dynamics of the non-Hermitian Hamiltonian
Figure 3: Left column: coloured chains illustrate how the logical qubits of two representative model Hamiltonians are mapped onto the physical Pegasus P16 topology of the D-Wave hardware. Panel (a) shows the native single-qubit problem $H_{1}$, Eq. (11), which embeds without chains. Panel (d) shows the non-native single-qubit problem $H_{3}$, Eq. (13), where minor embedding produces chains that span several unit cells. Middle column: ground-state success probability obtained on the two quantum annealers-Advantage2 (solid lines) and the earlier Advantage system (dashed lines)-as a function of the annealing time. Circles correspond to a temporal discretisation with two time points, triangles to three. Each marker averages 50005000 annealing runs; if no ground-state sample is observed the marker is omitted. Right column: expectation value $\langle\sigma_{z}\rangle$ reconstructed from the annealer output for the three lowest eigenstates of the target Hamiltonian. Blue triangles denote the ground state, green squares and red circles the first and second excited states, respectively. The black dashed curve is the exact unitary evolution computed with QuTiP [46, 47, 48, 49, 50] and serves as a baseline.
@@ III.2 Performance Scaling
Figure 4: Scaling properties of $\text{TTS}_{99}$ for Pegasus-native systems $(H_{1},H_{2},H_{5},H_{6},H_{7})$Equations˜11, 12, 15, 16 and 17, on the Advantage and Advantage2 QPU, as well as the classical solvers: Simulated Annealing (SA), both CPU and GPU based, and the GPU based VeloxQ. NN denotes the number of problem variables. Data points indicate values $\text{TTS}_{99}$, averaged over system size for all considered systems. Missing data points indicate failure to find the ground state. 5 (20) runs with 1000 samples each were performed on quantum (classical) solvers, for each system at time points $t\in\{2,3,\ldots,14\}$.
Figure 4 shows scaling of $\text{TTS}_{99}$ for different instances on quantum and classical solvers. Since the $\text{TTS}_{99}$ grows exponentially with the size of the system, we use the fitting function of the form
@@
where $D$ is the function’s offset and $N$ is the system size [53]. The most important fitting parameter is $\beta$, which can be related to average residual energy density, and subsequently loosely identified with an effective temperature of the generated ensemble of states [54]. It other words, it is a measure of computational complexity associated with a given problem type, and can be used as a quantitative way to compare different solvers, with larger value of $\beta$ indicating better performance. Of course, one should not expect this quantitiy to be universal across different QUBO problems/spin glass models, with evidence pointing to its dependence on interaction range, and spanning values from up to $\sim 10^{3}$ in case of Edwards-Anderson model on a cubic lattice [54], to as low as $\sim 4.5$ for all-to-all Sherrington-Kirkpatrick model [55]. In particular, it is not yet fully clear whether quantum solvers (e.g. nonadiabatic quantum annealing) exhibit any inherent advantage over fully classical one.
@@ IV Conclusion
For the set of dynamical problems tested, D-Wave Advantage2 produced the correct ground-state answer roughly ten times more often than its predecessor, and the increase held across all annealing schedules that we explored. In practical terms the probability of success rises while the time required to reach a fixed 99 % confidence level-often used as an application benchmark-grows far more slowly with problem size than on the earlier chip. These gains mirror the ongoing improvements in qubit coherence, control precision, and on-chip connectivity delivered by the latest fabrication cycle.
@@
The prospects for quantum hardware are encouraging. Our mapping requires additional qubits only in proportion to the number of simulated time steps, rather than the square or cubic growth familiar from many other encodings. As denser qubit lattices and wider coupler graphs move from prototype to production, medium-scale dynamical simulations covering tens to hundreds of effective degrees of freedom should become accessible within a single annealing run. Whether the next generation of processors merely matches or overtakes the best classical heuristics will depend on continued advances in both camps. The open-source benchmark suite released with this work provides a transparent reference against which that progress can be measured.

### v2

# Quantum-inspired dynamical models on quantum and classical annealers

@@ Abstract
We propose a practical, physics-inspired benchmarking suite to challenge both quantum and classical computers by mapping real-time quantum dynamics to a common optimization format. Using a parallel-in-time encoding, we convert the real-time propagator of an $n$-qubit, possibly non-Hermitian, Hamiltonian into quadratic unconstrained binary optimization (QUBO) instances that are executable in a solver-agnostic manner on quantum annealers and classical optimizers alike. This enables direct, like-for-like performance comparisons across fundamentally different computational paradigms. To stress-test the framework, we consider eight representative dynamical models spanning single-qubit rotations, multi-qubit entangling gates (Bell, GHZ, cluster), and PT-symmetric and other non-Hermitian generators, and evaluate success probability and time-to-solution as standard benchmarking metrics. Applying this methodology to two generations of D-Wave quantum annealers and to state-of-the-art classical solvers (Simulated Annealing and the GPU-accelerated VeloxQ), we find that Advantage2 consistently outperforms its predecessor, while VeloxQ retains the shortest absolute runtimes, reflecting the maturity of classical heuristics. We further extend the benchmarks to large-scale instances ($N\simeq 10^{5}$), establishing a demanding classical baseline for future hardware. Together, these results position the parallel-in-time QUBO framework as a versatile and physically motivated testbed for quantitatively tracking progress toward quantum-competitive simulation of dynamical systems.
@@ I Introduction
While our QUBO encoding builds on the parallel-in-time construction introduced in Ref. [27], the present work addresses a fundamentally different question. Rather than demonstrating feasibility on early-generation hardware, we establish a solver-agnostic, physics-derived benchmarking framework for quantum dynamics and use it to quantitatively assess the current performance gap between modern quantum annealers and state-of-the-art classical heuristics. In contrast to Ref. [27], we consider non-Hermitian and PT-symmetric generators, systematically distinguish native and non-native hardware embeddings, and benchmark two generations of D-Wave processors against GPU-accelerated classical solvers. Unlike Ref. [16], which focuses on synthetic or optimization-driven QUBO instances, our benchmarks are derived directly from continuous quantum dynamics, providing a physically motivated and application-relevant testbed for tracking hardware and algorithmic progress.
@@ II.1 Quantum Annealing
governed by QPU dependent anneal schedules $A(s)$ and $B(s)$ (cf. Fig. (LABEL:anneal_schedule)). If the evolution is sufficiently slow, according to the adiabatic theorem [28], the system remains in its instantaneous ground state, and the final state gives the solution to the original problem. In addition to standard forward annealing, we employ cyclic annealing as a refinement strategy for selected QUBO instances [29, 30, 31]. The key idea is to repeatedly “soften” the optimisation landscape around a candidate solution and then “re-freeze” it. Operationally, the problem Hamiltonian is kept fixed, while the annealing control parameter $s$ is driven through a short back-and-forth schedule: starting from a classical spin configuration at $s=1$ (problem-dominated regime), the schedule is partially reversed to an intermediate point $s_{\mathrm{p}}<1$ to temporarily restore the transverse-field contribution, and then returned to $s=1$ where the state is measured. By repeating this cycle, the solver can leave shallow local minima while still exploiting information contained in the current best configuration [32, 33, 34].
A single cycle can be described by a piecewise-linear control schedule $s(t)$,
$s(t)=\begin{cases}1-(1-s_{p})\dfrac{t}{\tau_{\downarrow}},&0\leq t<\tau_{\downarrow},\\[3.0pt] s_{p},&\tau_{\downarrow}\leq t<\tau_{\downarrow}+\tau_{p},\\[3.0pt] s_{p}+(1-s_{p})\dfrac{t-\tau_{\downarrow}-\tau_{p}}{\tau_{\uparrow}},&\begin{aligned} &\tau_{\downarrow}+\tau_{p}\leq t\\ &\leq\tau_{\downarrow}+\tau_{p}+\tau_{\uparrow},\end{aligned}\end{cases}$ (3)
where $\tau_{\downarrow}$ and $\tau_{\uparrow}$ set the reverse/forward ramp durations and $\tau_{\mathrm{p}}$ is an optional pause. The turnaround point $s_{\mathrm{p}}$ controls the locality of the search: larger $s_{\mathrm{p}}$ keeps the dynamics close to the input configuration (local refinement), whereas smaller $s_{\mathrm{p}}$ induces broader exploration. We execute $N_{\mathrm{cyc}}$ cycles and keep the lowest objective value observed over the entire cycle sequence. In practice, cyclic annealing often yields rapid early improvement followed by diminishing returns [cf. the plateau behaviour in Fig. (3)], so the cycle budget is best treated as a tunable trade-off between solution quality and runtime.
D-Wave’s superconducting quantum annealers are now applied across a spectrum of physics-driven tasks. Experiments on a 5000-qubit processor reproduced quantum critical dynamics in 3D spin glasses, extracting critical exponents beyond the reach of classical simulation [35]. In molecular science, an annealer-based eigensolver has recovered excited-state spectra for small molecules, showing that electronic-structure problems can be recast as Ising models amenable to quantum annealing [36]. Hybrid quantum-classical workflows have also tackled real-time traffic-flow optimisation, demonstrating measurable congestion reduction on realistic road networks [37, 38, 39, 40, 41, 42]. In computer science, an adiabatic implementation of Simon’s algorithm has been recently proposed and solved on D-Wave quantum annealers [43]. Furthermore, the D-Wave machine has been used as a platform for thermodynamic experiments to validate quantum fluctuation theorems and simulate quantum thermal machines [44, 45, 46, 47]. Progress notwithstanding, widespread deployment is limited by sparse qubit connectivity, which imposes costly minor-embeddings, and by analogue control errors and thermal noise that dilute prospective speed-ups [48, 49, 50, 51].
@@
Figure 3: Exemplary plateau behavior observed in cyclic annealing for instance $H_{3}$ with three time points. The blue line indicates the mean over the minimum energy found over each iteration, while the red dashed line shows the mean of the energies found by standard forward annealing. Shadowed box indicates the lower and upper quartile for all samples for forward annealing, and for lowest cycle energy for cyclic annealing. For this plot, we performed three cyclic annealing runs over 30 cycles and 30 forward annealing runs on Advantage2_system 1.10. The percentage in the box corresponds to the percentage with which lowest energy samples were found by each method.
@@ II.2 VeloxQ
VeloxQ is a fast and efficient, physics-inspired solver based on the theory of dynamical systems, designed for solving QUBO problems, which are crucial for various real-world optimisation tasks [52]. In a recent work, it has been extensively benchmarked against many state-of-the-art QUBO solvers, D-Wave quantum annealers among them [16]. Unlike them (or other physics-inspired optimisation approaches such as gate-based quantum computers), VeloxQ does not rely on technological advancements in hardware to reach its full potential, as it leverages conventional computing technology (such as GPUs). While it is a classical solver, it emulates the probabilistic nature of quantum annealers by producing a low-energy spectrum of solutions. By default, the solver operates with a carefully selected set of parameters that include instance-dependent tuning. However, it also offers flexibility, allowing users to adjust parameters to control the trade-off between solution quality and runtime based on problem knowledge or prior runs. Crucially, it offers stable and repeatable results, which do not suffer from unstable hardware. VeloxQ distinguishes itself from D-Wave quantum annealers in several significant ways. A key difference lies in their handling of problem topology: D-Wave’s quantum processing units (QPUs) have a fixed spin connectivity structure, and necessitate “minor embedding” for most real-world problems due to their specific topologies, which introduces substantial overhead. In contrast, VeloxQ natively handles all problem topologies without requiring such embedding. Furthermore, VeloxQ exhibits superior scalability, capable of tackling large-scale optimisation problems with up to $2\times 10^{8}$ sparsely connected variables, a feat currently intractable for D-Wave and its hybrid solutions, with estimates suggesting decades before quantum annealers could reach similar capabilities [16].
@@ II.4 Problem Instances
We examine entangled and non-entangled quantum systems, as well as non-Hermitian Hamiltonians that model quantum dynamics such as $\mathcal{P}\mathcal{T}$-symmetric qubits. To perform an accurate performance comparison between classical and quantum solvers, we select dynamical systems that can be precisely modelled as QUBO even with limited precision. Some of these dynamical systems produce QUBO instances that are native to the Pegasus graph, others require an embedding.
@@
Besides Hermitian operators, which are usually used as observables in quantum theory, quasi-Hermitian Hamiltonians allow to model unitary evolution as well. We simulate the behavior of a specific type of quasi-Hermitian system, the so-called $\mathcal{P}\mathcal{T}$-symmetric Hamiltonians. $\mathcal{P}\mathcal{T}$-symmetric systems remain invariant under the combined parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$) operator. Both of these transformations are hermitian and independent of each other, with $\mathcal{P},\mathcal{T}]=0 [$53]. Notably, systems with unbroken $\mathcal{P}\mathcal{T}$ symmetry have a real eigenspectrum and exhibit quasi-Hermitian behavior. A system is unbroken if the eigenstates of $H$ correspond to the eigenstates of the $\mathcal{P}\mathcal{T}$ operator. We evaluate the dynamics of the non-Hermitian Hamiltonian
Figure 4: Left column: coloured chains illustrate how the logical qubits of two representative model Hamiltonians are mapped onto the physical Pegasus P16 topology of the D-Wave hardware. Panel (a) shows the native single-qubit problem $H_{1}$, Eq. (12), which embeds without chains. Panel (d) shows the non-native single-qubit problem $H_{3}$, Eq. (14), where minor embedding produces chains that span several unit cells. Middle column: ground-state success probability obtained on the two quantum annealers-Advantage2 (solid lines) and the earlier Advantage system (dashed lines)-as a function of the annealing time. Circles correspond to a temporal discretisation with two time points, triangles to three. Each marker averages 50005000 annealing runs; if no ground-state sample is observed the marker is omitted. Right column: expectation value $\langle\sigma_{z}\rangle$ reconstructed from the annealer output of the three lowest energy states of the QUBO model of the target Hamiltonian. Blue triangles denote the ground state, green squares and red circles the first and second excited states, respectively. The black dashed curve is the exact unitary evolution computed with QuTiP [54, 55, 56, 57, 58] and serves as a baseline.
@@ III.2 Performance Scaling
Figure 5: Scaling properties of $\text{TTS}_{99}$ for Pegasus-native systems $(H_{1},H_{2},H_{5},H_{6},H_{7})$Equations˜12, 13, 16, 17 and 18, on the Advantage and Advantage2 QPU, as well as the classical solvers: Simulated Annealing (SA), both CPU and GPU based, and the GPU based VeloxQ. NN denotes the number of problem variables. Data points indicate values $\text{TTS}_{99}$, averaged over system size for all considered systems. Missing data points indicate failure to find the ground state. 5 (20) runs with 1000 samples each were performed on quantum (classical) solvers, for each system at time points $t\in\{2,3,\ldots,14\}$.
Figure 5 shows scaling of $\text{TTS}_{99}$ for different instances on quantum (using forward and cyclic annealing) and classical solvers. Since the $\text{TTS}_{99}$ grows exponentially with the size of the system, we use the fitting function of the form
@@
where $D$ is the function’s offset and $N$ is the system size [61]. The most important fitting parameter is $\beta$, which can be related to average residual energy density, and subsequently loosely identified with an effective temperature of the generated ensemble of states [62]. In other words, it is a measure of computational complexity associated with a given problem type, and can be used as a quantitative way to compare different solvers, with larger value of $\beta$ indicating better performance. Of course, one should not expect this quantitiy to be universal across different QUBO problems/spin glass models, with evidence pointing to its dependence on interaction range, and spanning values from up to $\sim 10^{3}$ in case of Edwards-Anderson model on a cubic lattice [62], to as low as $\sim 4.5$ for all-to-all Sherrington-Kirkpatrick model [63]. In particular, it is not yet fully clear whether quantum solvers (e.g. nonadiabatic quantum annealing) exhibit any inherent advantage over fully classical one.
Figure (5) shows that on the Advantage processor the fitted exponential scale parameter improves from $\beta\simeq 8.65$ (forward) to $\beta\simeq 9.82$ (cyclic), indicating that cycling can increase the likelihood of reaching the lowest-energy basin, but does not change the overall exponential character of the workload. This is consistent with the mechanism of cyclic annealing as a guided refinement step: it can help escape nearby suboptimal traps, yet it remains constrained by the same hardware noise floor and the same energy landscape structure [29].
Within the overlap region, the cyclic points show substantial scatter and do not exhibit a systematic reduction of $\mathrm{TTS}_{99}$ relative to forward annealing. A natural interpretation is that the extra control cycles add runtime per attempt, and—once the best reachable energy has been found—additional cycles mainly accumulate overhead without proportionally increasing the probability of hitting the true ground state. This interpretation is directly supported by Fig. (3), where the best-found objective value drops quickly in the first few cycles and then plateaus.
@@
Finally, we extend our benchmarks beyond instances that fit (even with embedding) on current generation annealers. This allows us to establish a classical baseline, within the framework introduced in this work, against which the performance of future quantum devices might be measured. We select three Pegasus-native systems $(H_{1},H_{2},H_{6})$, and construct corresponding QUBO instances with number of discretization timepoints $\{100,1000,2000,5000,10000\}$ (systems $H_{1},H_{2}$) and $\{100,1000,2000\}$ (system $H_{6}$). This results in QUBO instances with sizes ranging from $N=800$ to $N=80000$ variables. Since it is infeasible for a heuristic solvers to obtain the true ground state in reasonable time, we relax our performance gauge and use the so-called time-to-epsilon $\mathrm{TT}\varepsilon$ [12, 15], which measures the expected time needed to reach a solution with optimality gap at most $\varepsilon$,
$\mathrm{TT}\varepsilon=\frac{\log(1-0.99)}{\log(1-P_{E\leq E_{0}+\varepsilon|E_{0}|})}T_{r}.$ (22)
For the comparison, we consider only the GPU-based VeloxQ solver, and the GPU version of Simulated Annealing. To estimate the success probabiltiy, we performed $100$ independent runs in the former case, and $25$ in the latter. Obtained values of $\mathrm{TT}\varepsilon$ were then optimized over the relevant control parameters of the solvers, them being the number of steps and simultaneously evolved trajectories for VeloxQ, and number of Monte Carlo Sweeps and simultaneously evolved replicas for SA. The results are shown in Figure (6). Both solvers exhibit power-law like scaling, with slope increasing with decreasing $\varepsilon$, and eventually diverging to infinity as $\varepsilon\to 0$. This indicates the transition to the exponential scaling of time-to-solution and is consistent with results presented in Figure 21. We see that for all considered optimality targets $\varepsilon\in\{1.0,2.0,3.5,5.0\}\%$, VeloxQ produces solutions of desired quality significantly faster (on average), than Simulated Annealing. The only exception is the case of strictest target of $\varepsilon=1\%$, where both solvers cannot reach it for the largest instances, but VeloxQ fails already for smaller systems than SA. Nevertheless, VeloxQ proves itself as a capable solver, especially in the realm of approximate optimization of very large problems.
Figure 6: Scaling of time-to-epsilon $\mathrm{TT}\varepsilon$ for GPU-based classical heuristics, VeloxQ and SA, on large-scale QUBO instances (NN up to $10^{5}$), generated from systems $H_{1},H_{2},H_{6}$. Both algorithms display characteristic power-law scaling, where the steepness of the slope is inversely related to the target optimality gap $\varepsilon$. This behavior indicates a crossover toward exponential complexity as the target error vanishes ($\varepsilon\to 0$). For almost all approximate targets shown ($\varepsilon\in\{1.0,2.0,3.5,5.0\}\%$), VeloxQ demonstrates a consistent advantage over Simulated Annealing. The only exception is for $\varepsilon=1.0\%$, where VeloxQ fails to reach the target already for smaller instances than SA.
@@ IV Conclusion
For the set of dynamical problems tested, D-Wave Advantage2 produced the correct ground-state answer roughly ten times more often than its predecessor, and the increase held across all annealing schedules that we explored. In practical terms the probability of success rises while the time required to reach a fixed 99 % confidence level-often used as an application benchmark-grows far more slowly with problem size than on the earlier chip. These gains mirror the ongoing improvements in qubit coherence, control precision, and on-chip connectivity delivered by the latest fabrication cycle. We additionally tested cyclic annealing as a refinement step for selected instances and found that it typically improves the best objective rapidly in the first few cycles before saturating. At the aggregate level, this translates into only modest changes in time-to-solution scaling. These observations indicate that cyclic annealing is most useful as a targeted local-improvement heuristic, while the dominant performance gains in our benchmark are driven by the hardware generation upgrade to Advantage2.
@@
The prospects for quantum hardware are encouraging. Our mapping requires additional qubits only in proportion to the number of simulated time steps, rather than the square or cubic growth familiar from many other encodings. As denser qubit lattices and wider coupler graphs move from prototype to production, medium-scale dynamical simulations covering tens to hundreds of effective degrees of freedom should become accessible within a single annealing run. Whether the next generation of processors merely matches or overtakes the best classical heuristics will depend on continued advances in both camps.
Our large-scale benchmarks ($N$ up to $\sim 10^{5}$) confirm that VeloxQ significantly outperforms Simulated Annealing, exhibiting favorable power-law scaling. This establishes a high-performance classical baseline that helps to indentify the threshold for future quantum advantage. The open-source benchmark suite released with this work provides a transparent reference against which that progress can be measured.